# T-JEPA: Augmentation-Free Self-Supervised Learning for Tabular Data

### v1

@@ 1 Introduction
Most self-supervised approaches include deep models, which have excelled in applications that include images or text. However, using neural networks for tabular data still remains challenging (Shwartz-Ziv and Armon, 2021). Indeed, Grinsztajn et al. (2022) discuss how the inherent heterogeneity of tabular data makes learning from this data structure using neural networks difficult. Nonetheless, recent work has investigated finding effective training procedures and novel architectures to learn from tabular data using neural networks. Recent advancement include improved training procedures (Kadra et al., 2021; Gorishniy et al., 2021b; Hollmann et al., 2023), representation learning for tabular data (Ye et al., 2024; Bahri et al., 2022; Zhu et al., 2023) or novel architectures (Somepalli et al., 2021; Kossen et al., 2021). Despite these recent advancements, leveraging self-supervised learning for tabular data remains strenuous as most methods involve data augmentations to construct multiple views of the same data sample. While augmentations can be relatively straightforward for images or text data, constructing meaningful augmentation for tabular data is non-trivial. Augmentations for image samples often include cropping, rotation, or color alteration, while for text samples, this can include token masking or token replacement. These corruptions or augmentations of samples are domain-specific and hard to translate for structured tabular data as they can generate samples outside the data manifold.
@@
Figure 1: T-JEPA training pipeline. In step (a) a sample $\mathbf{x}\in\mathbb{R}^{d}$ is pre-processed and masked as detailed in equation 1 and fed to the context encoder to obtain a representation in $\mathbb{R}^{l_{\mathbf{m}}\times h}$ where $l_{\mathbf{m}}$ is the number of unmasked features for context mask $\mathbf{m}$. In step (b) the unmasked representation of sample $\mathbf{x}$ is fed to the target encoder and the features’ representations are selected according to the corresponding target masks, as shown in equation 4. In step (c) the output of the context encoder is fed to the predictor to obtain a prediction for each target mask used in step (b). In step (d) we compute the $\ell_{2}$-distance between the target representations and their predictions.
As discussed by Assran et al. (2023), self-supervised learning for representation learning includes three types of approaches. First, Joint-Embedding Architecture (JEA) usually involves two encoders that learn to output similar embeddings for similar samples while ensuring distant embeddings for dissimilar samples. Second, Generative Architecture that aims at reconstructing a sample from a corrupted version of this sample, e.g., mask reconstruction. Third, Joint-Embedding Predictive Architecture (JEPA) resembles Generative Architecture as it relies on a similar task but in the latent space rather than the original data space. JEPA-based method consist in predicting a sample’s representation in the embedding space from the embedded representation of a corrupted version of this sample. Recently, Assran et al. (2023) have proposed I-JEPA, a novel self-supervised approach targeted for images that does not involve augmentations. Their approach used for pre-training offered significant improvement in classification tasks on images. Following their path, other works have extended their approach to video (Bardes et al., 2024), audio (Fei et al., 2024), and graphs (Skenderi et al., 2023). The present work aims to adapt JEPA for tabular data as a pre-training model to foster performance on classification and regression tasks. Recent work (Thimonier et al., 2024; Kossen et al., 2021) has demonstrated that mask reconstruction can be a relevant pretext task for representation learning of tabular data. This work extends this mask reconstruction paradigm from the data space to the latent space. Adapting such approach to structured data is particularly relevant as it avoids constructing ad-hoc data augmentations that are challenging to construct for this data type.
@@ 2 Related Work
Representation learning for tabular data has caught increasing attention in recent years. Gorishniy et al. (2021a) extensively investigate the benefits of pre-training models on tabular data to enhance performance. In other work, Somepalli et al. (2021); Kossen et al. (2021) propose a pretraining procedure to foster the performance of their novel transformer-based architectures. Parallel to that, some works have focused entirely on proposing self-supervised methods for representation learning of tabular data. One of the first approaches, VIME (Yoon et al., 2020), proposes to augment the existing reconstruction task with estimating mask vectors from corrupted tabular data. Bahri et al. (2022) propose Scarf, a simple method based on contrastive learning in which different views of a sample are obtained by corrupting a random subset of features. Recent work has also investigated prototype-based representation learning for tabular data as PTaRL (Ye et al., 2024). Parallel to this, Ucar et al. (2021) propose SubTab that divides the input features to multiple subsets and to perform a pretext task close to mask reconstruction. Other works, such as XTab (Zhu et al., 2023), TransTab (Wang and Sun, 2022) or UniTabE (Yang et al., 2024), propose self-supervised representation learning for cross-table pretraining. Wu et al. (2024) discuss the concepts of salient and mutual information and emphasize their key role in producing meaningful sample representations. They propose SwitchTab, which aims to foster the decoupling between the salient and mutual information contained in a sample to produce its representations. Finally, Lee et al. (2024) emphasize the necessity of correctly handling the heterogeneous features of tabular data. Somewhat close to our approach, their method consists in binning the values of each feature and proceeds to use as a pretext task the reconstruction of the bin indices rather than the value in the original feature space.
@@ 3 Method
Each of the $d$ features is equipped with a learned linear layer, $\texttt{Linear}(e_{j},h),\forall j\in\{1,\dots,d\}$, that embeds the $e_{j}$-dimensional representation into an $h$-dimensional space. We pass each of the $l_{\mathbf{m}}$ unmasked features’ encoded representations through their corresponding linear layers. We also learn $h$-dimensional index and feature-type embeddings. Both are added to the embedded representation of sample $\mathbf{x}$. Let $\mathbf{z}_{\mathbf{x}}^{\mathbf{m}}\in\mathbb{R}^{l_{\mathbf{m}}\times h}$ denote the obtained embedded representation of sample $\mathbf{x}$ with mask $\mathbf{m}$.
@@
We also include a regularizing token [REG] inspired from the register token first proposed in (Darcet et al., 2024) for ViT’s. We append this token to the obtained $\tilde{\mathbf{E}}(\mathbf{x})$ representation displayed in equation 1. This token is also equipped with a learned embedding layer. This token is only used to train T-JEPA and is discarded when training supervised classifiers on the downstream task. We later discuss the necessity of including such token in section 5.2 and observe that it acts as a regularizing method to escape training regimes leading to representation collapse. For simplicity, we do not explicitly include the regularizing token in the rest of the method description hereafter.
@@
The masking strategy differs between the context and target encoders. We sample several masks for each sample. Context masks are used to mask the samples before feeding them to the embedding module and context encoder. On the contrary, the target masks are used to construct the target representation after passing them through the embedding module and target encoder. Note that in both context and target masking, the regularizing token is never masked. Hence, the input of the context encoder is a masked representation of a sample $\mathbf{x}$, $\mathbf{z}_{\mathbf{x}}^{\mathbf{m}}$. In contrast, the target encoder receives as an input the embedded representation $\mathbf{z}_{\mathbf{x}}^{\mathbf{0}_{d}}$, where $\mathbf{0}_{d}$ is the $d$-dimensional null vector. Then, the target mask is used to mask the corresponding encoded feature representations of $\mathbf{z}_{\mathbf{x}}^{\mathbf{0}_{d}}$ as shown in equation 4. For both context and target encoders, we set a minimum and maximum share of features to be masked simultaneously and randomly sample a share in that interval. We construct context and target masks so that there is no overlap between context and target masks but allow overlap within context masks and target masks. Let $M_{\text{context}},M_{\text{target}}$ denote the set of sampled context and target masks, respectively.
@@ 4.1 Experimental Setting
Following previous work (Ye et al., 2024), we experiments on $6$ datasets with heterogeneous features to test the effectiveness of T-JEPA. We test our approach on several supervised tabular deep learning tasks such as binary and multi-class classification, as well as regression. We use as performance metrics Accuracy ($\uparrow$) and RMSE ($\downarrow$) for classification and regression respectively. The datasets we include in our experiments are Adult (AD) (Kohavi et al., 1996), Higgs (HI) (Vanschoren et al., 2014), Helena (HE) (Guyon et al., 2019), Jannis (JA) (Guyon et al., 2019), ALOI (AL) (Geusebroek et al., 2005) and California housing (CA) (Pace and Barry, 1997). We summarize the characteristics of all $6$ datasets in Table 5 in appendix B.
@@
To assess whether our self-supervised approach can foster performance on tabular applications, we compare the performance of several widely used tabular approaches with and without our self-supervised pre-training. We test our method on MLP (Taud and Mas, 2018), DCNV2 (Wang et al., 2021), ResNet (He et al., 2016), AutoInt (Song et al., 2019) and FT-Transformer (Gorishniy et al., 2021b). For completeness, we also compare the performance on the downstream task to other SSL methods like PTaRL (Ye et al., 2024), SwitchTab (Wu et al., 2024), BinRecon (Lee et al., 2024) and VIME (Yoon et al., 2020) (see Table 6 in Appendix E). Similarly, as often considered as the go-to methods for supervised tasks on tabular data, we also compare the performance of T-JEPA to XGBoost (Chen and Guestrin, 2016) and CatBoost (Prokhorenkova et al., 2018).
@@
We split each dataset into training/validation/test sets (80/10/10) which were used for selecting both the hyperparameters of T-JEPA and of the models used for the downstream task. More precisely, in a model-agnostic manner, we relied on a systematic approach to train and evaluate the embedding space generated by T-JEPA. First, following previous work (Assran et al., 2023; Bardes et al., 2022) T-JEPA was trained on the training set and we conducted a hyperparameter tuning using a linear probe on the validation set to select the best configuration for each dataset. Second, we used the trained context encoder to generate data representations on which the subsequent model were trained. We refer the reader to appendix A.2 for more details on hyperparameter selection. Every experiment can be reproduced with the code made available on GitHub.
@@
To adapt the T-JEPA representations to each model’s input dimensions, we added a projection layer. We experimented with several techniques, including linear flattening, linear per-feature transformation, convolutional projections, and max and mean pooling. We refer the reader to appendix A.3 for more details. The projection layer was jointly trained with the downstream task and tailored to each model. Hyperparameter tuning was performed based on validation set performance, and final evaluations were conducted on the test set, comparing results to models trained on the original dataset representations. To ensure the robustness and reliability of our findings, we conducted the experiment 20 times. We report the mean and standard deviation of the performance metrics.
Table 1: Performance metrics for different models with and without T-JEPA representation across various datasets. Metrics for the models without T-JEPA are averaged over $10$ runs and taken from Ye et al. (2024). Regarding models augmented by T-JEPA, we report an average over 2020 runs, see Table 6 in appendix E for comparison with existing methods. We report in bold the metric that wins between the raw data representation and T-JEPA’s. We report in blue the best performing deep method and underline the overall best metric for a dataset.
AD $\uparrow$ HE $\uparrow$ JA $\uparrow$ AL $\uparrow$ CA $\downarrow$ HI $\uparrow$ Wins
Baseline Neural Networks
MLP $0.825$ $0.352$ $0.672$ $0.917$ $0.518$ $\mathbf{0.681}$ 11
+T-JEPA $\mathbf{0.866}$ $\mathbf{0.400}$ $\underline{\mathbf{0.728}}$ $\mathbf{0.961}$ $\mathbf{0.468}$ $0.517$ $\mathbf{5}$
DCNv2 $0.826$ $0.34$ $0.662$ $0.905$ $0.502$ $\mathbf{0.681}$ 11
+T-JEPA $\mathbf{0.861}$ $\mathbf{0.399}$ $\mathbf{0.723}$ $\mathbf{0.955}$ $\underline{\mathbf{0.420}}$ $0.525$ $\mathbf{5}$
ResNet $0.813$ $0.354$ $0.666$ $0.919$ $0.537$ $0.682$ 0
+T-JEPA $\mathbf{0.865}$ $\underline{\mathbf{0.401}}$ $\mathbf{0.718}$ $\underline{\mathbf{0.964}}$ $\mathbf{0.441}$ $\mathbf{0.705}$ $\mathbf{6}$
AutoInt $0.823$ $0.338$ $0.653$ $0.894$ $0.507$ $\mathbf{0.685}$ 11
+T-JEPA $\mathbf{0.866}$ $\mathbf{0.351}$ $\mathbf{0.710}$ $\mathbf{0.938}$ $\mathbf{0.448}$ $0.517$ $\mathbf{5}$
FT-Trans $0.827$ $0.363$ $0.675$ $0.913$ $0.486$ $\mathbf{0.689}$ 11
+T-JEPA $\mathbf{0.864}$ $\mathbf{0.384}$ $\mathbf{0.708}$ $\mathbf{0.921}$ $\mathbf{0.444}$ $0.551$ $\mathbf{5}$
Gradient Boosted Decision Trees (GBDT)
XGBoost $\underline{0.874}$ $0.368$ $0.720$ $0.951$ $0.462$ $\underline{0.729}$ N/A
CatBoost $0.873$ $0.381$ $0.721$ $0.946$ $0.430$ $0.726$ N/A

### v3

@@ 1 Introduction
Most self-supervised approaches include deep models, which have excelled in applications that include images or text. However, using neural networks for tabular data still remains challenging (Shwartz-Ziv and Armon, 2021). Indeed, Grinsztajn et al. (2022) discuss how the inherent heterogeneity of tabular data makes learning from this data structure using neural networks difficult. Nevertheless, recent work has investigated finding effective training procedures and novel architectures to learn from tabular data using neural networks. Recent advances include improved training procedures (Kadra et al., 2021; Gorishniy et al., 2021; Hollmann et al., 2023), representation learning for tabular data (Ye et al., 2024; Bahri et al., 2022; Zhu et al., 2023) or novel architectures (Somepalli et al., 2021; Kossen et al., 2021). Despite these recent advances, leveraging self-supervised learning for tabular data remains strenuous as most methods involve data augmentations to construct multiple views of the same data sample. While augmentations can be relatively straightforward for images or text data, constructing meaningful augmentation for tabular data is non-trivial. Augmentations for image samples often include cropping, rotation, or color alteration, while for text samples, this can include token masking or token replacement. These corruptions or augmentations of samples are domain-specific and hard to translate for structured tabular data as they can generate samples outside the data manifold.
@@
Figure 1: T-JEPA training pipeline. In step (a) a sample $\mathbf{x}\in\mathbb{R}^{d}$ is pre-processed and masked as detailed in equation 1 and fed to the context encoder to obtain a representation in $\mathbb{R}^{l_{\mathbf{m}}\times h}$ where $l_{\mathbf{m}}$ is the number of unmasked features for context mask $\mathbf{m}$. In step (b) the unmasked representation of sample $\mathbf{x}$ is fed to the target encoder and the features’ representations are selected according to the corresponding target masks, as shown in equation 4 and Figure 7. In step (c) the output of the context encoder is fed to the predictor to obtain a prediction for each target mask used in step (b). In step (d) we compute the $\ell_{2}$-distance between the target representations and their predictions.
As discussed by Assran et al. (2023), self-supervised learning for representation learning includes three types of approaches. First, Joint-Embedding Architecture (JEA) usually involves two encoders that learn to output similar embeddings for similar samples while ensuring distant embeddings for dissimilar samples. Second, Generative Architecture that aims at reconstructing a sample from a corrupted version of this sample, e.g., mask reconstruction. Third, Joint-Embedding Predictive Architecture (JEPA) resembles Generative Architecture as it relies on a similar task but in the latent space rather than the original data space. JEPA-based method consist in predicting a sample’s representation in the embedding space from the embedded representation of a corrupted version of this sample. Recently, Assran et al. (2023) have proposed I-JEPA, a novel self-supervised approach targeted for images that does not involve augmentations. Their approach used for pre-training offered significant improvement in classification tasks on images. Following their path, other works have extended their approach to video (Bardes et al., 2024), audio (Fei et al., 2024), and graphs (Skenderi et al., 2023). The present work aims to adapt JEPA for tabular data as a pre-training model to foster performance on classification and regression tasks. Recent work (Kossen et al., 2021; Ucar et al., 2021; Thimonier et al., 2024) has demonstrated that mask reconstruction can be a relevant pretext task for representation learning of tabular data. This work extends this mask reconstruction paradigm from the data space to the latent space. Adapting such approach to structured data is particularly relevant as it avoids constructing ad-hoc data augmentations that are challenging to construct for this data type.
@@ 2 Related Work
Representation learning for tabular data has caught increasing attention in recent years. Gorishniy et al. (2021) extensively investigate the benefits of pre-training models on tabular data to enhance performance. In other works, Somepalli et al. (2021) and Kossen et al. (2021) propose a pretraining procedure to foster the performance of their novel transformer-based architectures for tabular data. Parallel to that, some works have focused entirely on proposing self-supervised methods for representation learning of tabular data. One of the first approaches, VIME (Yoon et al., 2020), proposes to augment the existing reconstruction task with estimating mask vectors from corrupted tabular data. Bahri et al. (2022) propose Scarf, a simple method based on contrastive learning in which different views of a sample are obtained by corrupting a random subset of features. Recent work has also investigated prototype-based representation learning for tabular data such as PTaRL (Ye et al., 2024). Other works, such as XTab (Zhu et al., 2023), TransTab (Wang and Sun, 2022) or UniTabE (Yang et al., 2024), propose self-supervised representation learning for cross-table pretraining. Wu et al. (2024) discuss the concepts of salient and mutual information and emphasize their key role in producing meaningful sample representations. They propose SwitchTab, which aims to foster the decoupling between the salient and mutual information contained in a sample to produce its representations. Lee et al. (2024) emphasize the necessity of correctly handling the heterogeneous features of tabular data. Somewhat close to our approach, their method consists in binning the values of each feature and proceeds to use as a pretext task the reconstruction of the bin indices rather than the value in the original feature space. Also related to our approach, Sui et al. (2024) propose an augmentation-free method that can be applied to tabular data. Their method consists in simultaneously reconstructing multiple randomly generated data projection functions to generate a data representation. Finally, most related to our method, Ucar et al. (2021) propose SubTab that divides the input features to multiple subsets to perform a pretext task close to mask reconstruction. The core difference with T-JEPA lies in the fact SubTab performs mask reconstruction in the original dataspace while T-JEPA performs this task in the embedded space. We provide in Appendix F.2 a more comprehensive description of each SSL methods relevant to the present work.
@@ 3 Method
Each of the $d$ features is equipped with a learned linear layer, $\texttt{Linear}(e_{j},h),\forall j\in\{1,\dots,d\}$, that embeds the $e_{j}$-dimensional representation into an $h$-dimensional space. We pass each of the $l_{\mathbf{m}}$ unmasked features’ encoded representations through their corresponding linear layers. We also learn $h$-dimensional index and feature-type embeddings following standard practice when leveraging transformers for tabular data. Both are added to the embedded representation of sample $\mathbf{x}$. Let $\mathbf{z}_{\mathbf{x}}^{\mathbf{m}}\in\mathbb{R}^{l_{\mathbf{m}}\times h}$ denote the obtained embedded representation of sample $\mathbf{x}$ with mask $\mathbf{m}$.
@@
We also include a regularizing token [REG] inspired from the register token first proposed in (Darcet et al., 2024) for ViT’s. We append this token to the obtained $\tilde{\mathbf{E}}(\mathbf{x})$ representation displayed in equation 1. This token is also equipped with a learned embedding layer. This token is only used to train T-JEPA and is discarded when training supervised classifiers on the downstream task. See Fig 7 in Appendix H for an illustration. We later discuss the necessity of including such token in section 5.2 and observe that it acts as a regularizing method to escape training regimes leading to representation collapse. For simplicity, we do not explicitly include the regularizing token in the rest of the method description hereafter. See Appendix H for a more detailed discussion on regularization tokens.
@@
The masking strategy differs between the context and target encoders. We sample several masks for each sample. Context masks are used to mask the samples before feeding them to the embedding module and context encoder. On the contrary, the target masks are used to construct the target representation after passing them through the embedding module and target encoder. Note that in both context and target masking, the regularizing token is never masked. Hence, the input of the context encoder is a masked representation of a sample $\mathbf{x}$, $\mathbf{z}_{\mathbf{x}}^{\mathbf{m}}$. In contrast, the target encoder receives as an input the embedded representation $\mathbf{z}_{\mathbf{x}}^{\mathbf{0}_{d}}$, where $\mathbf{0}_{d}$ is the $d$-dimensional null vector. Then, the target mask is used to mask the corresponding encoded feature representations of $\mathbf{z}_{\mathbf{x}}^{\mathbf{0}_{d}}$ as shown in equation 4. For both context and target encoders, we set a minimum and maximum share of features to be masked simultaneously and randomly sample a share in that interval. Let $M_{\text{context}},M_{\text{target}}$ denote the set of sampled context and target masks, respectively. We construct $M_{\text{context}},M_{\text{target}}$ such that intra-overlaps are permitted (masks from the same set can overlap), but inter-overlaps are not permitted (a mask from $M_{\text{context}}$ cannot overlap with a mask from $M_{\text{target}}$).
@@ 4.1 Experimental Setting
Following previous work (Ye et al., 2024), we experiments on $7$ datasets with heterogeneous features to test the effectiveness of T-JEPA. We test our approach on several supervised tabular deep learning tasks such as binary and multi-class classification, as well as regression. We use as performance metrics Accuracy ($\uparrow$) and RMSE ($\downarrow$) for classification and regression respectively. The datasets we include in our experiments are Adult (AD) (Kohavi et al., 1996), Higgs (HI) (Vanschoren et al., 2014), Helena (HE) (Guyon et al., 2019), Jannis (JA) (Guyon et al., 2019), ALOI (AL) (Geusebroek et al., 2005) and California housing (CA) (Pace and Barry, 1997). We also add MNIST (interpreted as a tabular data) to our benchmark following Yoon et al. (2020). We summarize the characteristics of all $7$ datasets in Table 11 in appendix C.
@@
To assess whether our self-supervised approach can foster performance on tabular applications, we compare the performance of several widely used tabular approaches with and without our self-supervised pre-training. We test our method on MLP (Taud and Mas, 2018), DCNV2 (Wang et al., 2021a), ResNet (He et al., 2016), AutoInt (Song et al., 2019) and FT-Transformer (Gorishniy et al., 2021). Similarly, as often considered as the go-to methods for supervised tasks on tabular data, we also compare the performance of T-JEPA augmented models to XGBoost (Chen and Guestrin, 2016) and CatBoost (Prokhorenkova et al., 2018). To further assess the relevance of T-JEPA to foster performance on supervised tasks, we compare the performance of MLP and ResNet when trained on the original dataspace representations and those generated by PTaRL (Ye et al., 2024), SwitchTab (Wu et al., 2024), BinRecon (Lee et al., 2024), SubTab (Ucar et al., 2021) VIME (Yoon et al., 2020) and T-JEPA.
Table 1: Performance metrics for different downstream models trained on the original dataspace and the generated T-JEPA representation across datasets. We also include for completeness the performance of XGBoost (Chen and Guestrin, 2016) and CatBoost (Prokhorenkova et al., 2018) as a baseline. We report an average over 2020 runs and the corresponding standard deviation. We report in bold the metric that wins between the raw data representation and the augmented representations. We underline the overall best metric for a dataset.
AD $\uparrow$ HE $\uparrow$ JA $\uparrow$ AL $\uparrow$ CA $\downarrow$ HI $\uparrow$ MNIST $\uparrow$ Baseline Neural Networks MLP $0.827$ $\pm 1e^{-3}$ $0.353$ $\pm 1e^{-3}$ $0.672$ $\pm 1e^{-3}$ $0.916$ $\pm 4e^{-3}$ $0.511$ $\pm 3e^{-3}$ $\mathbf{0.681}$ $\pm 4e^{-3}$ $0.978$ $\pm 2e^{-3}$ +T-JEPA $\mathbf{0.866}$ $\pm 2e^{-3}$ $\mathbf{0.400}$ $\pm 4e^{-3}$ $\underline{\mathbf{0.728}}$ $\pm 3e^{-3}$ $\mathbf{0.961}$ $\pm 6e^{-3}$ $\mathbf{0.468}$ $\pm 4e^{-2}$ $0.517$ $\pm 8e^{-2}$ $\mathbf{0.983}$ $\pm 1e^{-3}$ DCNv2 $0.829$ $\pm 4e^{-3}$ $0.347$ $\pm 3e^{-3}$ $0.662$ $\pm 3e^{-3}$ $0.905$ $\pm 3e^{-3}$ $0.504$ $\pm 4e^{-3}$ $\mathbf{0.683}$ $\pm 3e^{-3}$ $0.971$ $\pm 1e^{-3}$ +T-JEPA $\mathbf{0.861}$ $\pm 2e^{-3}$ $\mathbf{0.399}$ $\pm 3e^{-3}$ $\mathbf{0.723}$ $\pm 2e^{-3}$ $\mathbf{0.955}$ $\pm 2e^{-3}$ $\mathbf{0.420}$ $\pm 3e^{-2}$ $0.525$ $\pm 8e^{-2}$ $\mathbf{0.981}$ $\pm 2e^{-3}$ ResNet $0.814$ $\pm 7e^{-3}$ $0.351$ $\pm 2e^{-3}$ $0.666$ $\pm 3e^{-3}$ $0.919$ $\pm 2e^{-3}$ $0.534$ $\pm 2e^{-3}$ $0.674$ $\pm 4e^{-3}$ $0.979$ $\pm 1e^{-3}$ +T-JEPA $\mathbf{0.865}$ $\pm 3e^{-3}$ $\mathbf{0.401}$ $\pm 2e^{-3}$ $\mathbf{0.718}$ $\pm 3e^{-3}$ $\mathbf{\underline{0.964}}$} $\pm 1e^{-3}$ $\mathbf{0.441}$ $\pm 8e^{-2}$ $\mathbf{0.705}$ $\pm 5e^{-3}$ $\mathbf{0.983}$ $\pm 2e^{-3}$ AutoInt $0.823$ $\pm 1e^{-3}$ $0.338$ $\pm 3e^{-3}$ $0.653$ $\pm 6e^{-3}$ $0.894$ $\pm 2e^{-3}$ $0.501$ $\pm 3e^{-3}$ $\mathbf{0.694}$ $\pm 3e^{-3}$ $0.901$ $\pm 6e^{-3}$ +T-JEPA $\mathbf{0.866}$ $\pm 2e^{-3}$ $\mathbf{0.351}$ $\pm 3e^{-3}$ $\mathbf{0.710}$ $\pm 3e^{-3}$ $\mathbf{0.938}$ $\pm 4e^{-3}$ $\mathbf{0.448}$ $\pm 2e^{-2}$ $0.517$ $\pm 8e^{-2}$ $\mathbf{0.978}$ $\pm 1e^{-3}$ FT-Trans $0.821$ $\pm 7e^{-3}$ $0.363$ $\pm 2e^{-3}$ $0.677$ $\pm 2e^{-3}$ $0.913$ $\pm 3e^{-3}$ $0.473$ $\pm 5e^{-3}$ $\mathbf{0.684}$ $\pm 5e^{-3}$ $0.811$ $\pm 5e^{-2}$ +T-JEPA $\mathbf{0.864}$ $\pm 1e^{-3}$ $\mathbf{0.384}$ $\pm 5e^{-3}$ $\mathbf{0.708}$ $\pm 5e^{-3}$ $\mathbf{0.921}$ $\pm 1e^{-2}$ $\mathbf{0.444}$ $\pm 1e^{-1}$ $0.551$ $\pm 6e^{-2}$ $\mathbf{0.966}$ $\pm 2e^{-3}$ Gradient Boosted Decision Trees (GBDT) XGBoost $\underline{0.872}$ $\pm 5e^{-4}$ $0.375$ $\pm 1.2e^{-3}$ $0.721$ $\pm 1e^{-3}$ $0.951$ $\pm 1e^{-3}$ $0.433$ $\pm 2e^{-3}$ $\underline{0.729}$ $\pm 1e^{-3}$ $0.980$ $\pm 1e^{-3}$ CatBoost $0.873$ $\pm 1e^{-3}$ $0.381$ $\pm 1e^{-3}$ $0.721$ $\pm 1e^{-3}$ $0.946$ $\pm 9e^{-4}$ $0.430$ $\pm 7e^{-4}$ $0.726$ $\pm 8e^{-4}$ $0.972$ $\pm 3e^{-3}$
@@
We split each dataset into training/validation/test sets (80/10/10) which were used for selecting both the hyperparameters of T-JEPA and of the models used for the downstream task. More precisely, in a model-agnostic manner, we relied on a systematic approach to train and evaluate the embedding space generated by T-JEPA. First, following previous work (Assran et al., 2023; Bardes et al., 2022) T-JEPA was trained on the training set and we conducted a hyperparameter tuning using a linear probe on the validation set to select the best configuration for each dataset. Second, we used the trained context encoder to generate data representations on which the subsequent model were trained. We refer the reader to appendix A.2 for more details on hyperparameter selection. Every experiment can be reproduced with the code provided in the supplementary material.
@@
To adapt the T-JEPA representations to each model’s input dimensions, we added a projection layer. We experimented with several techniques, including linear flattening, linear per-feature transformation, convolutional projections, and max and mean pooling. We refer the reader to appendix A.3 for more details. The projection layer was jointly trained with the downstream task and tailored to each model. Hyperparameter tuning was performed based on validation set performance, and final evaluations were conducted on the test set, comparing results to models trained on the original dataset representations. To ensure the robustness and reliability of our findings, we conducted the experiment 20 times. We report the mean and standard deviation of the performance metrics. We refer the reader to appendix B for details on the downstream models’ hyperparameters.